# From Fibers to Cells: Fourier-Based Registration Enables Virtual Cresyl Violet Staining From 3D Polarized Light Imaging

### v1

@@ 2.1.2 Tissue shrinkage correction
Deformations induced by histological processing include a volumetric shrinkage of brain tissue Amunts et al., (2005). To correct this effect, the extent of shrinkage is usually estimated from the ratio between the histological processed and true brain volume represented by the fresh weight of the whole brain Amunts et al., (2005, 2013) or an MRI reference Wagstyl et al., (2020). To estimate the shrinkage of the 3D-PLI acquisition, a linear registration of the 3D reconstructed blockface volume and the postmortem MRI of the same brain was performed. The determinant of the transformation matrix showed a global volume change below 1%, which can be neglected. In a second step, we calculated 2D shrinkage factors for each brain section considered in this work as the quotient between the area occupied by tissue in the 3D-PLI measurement and its corresponding blockface image. For test section #559, we estimated a 2D shrinkage factor of 0.97 meaning a slight expansion of the 3D-PLI measurement relative to its original area within the MRI. This observation is consistent throughout all sections with 2D shrinkage factors between 0.95 and 0.99. We report in-plane cell sizes measured with 3D-PLI multiplied by an individual 2D shrinkage factor per section, assuming a uniform area change of cells and surrounding tissue.
@@ 2.2 Initial cross-modality alignment
Performing the initial affine registration reveals remaining nonlinear deformations as shown in Fig. 2D. Since nonlinear deformations typically have low spatial frequencies, causing smooth, large-scale distortions, we expected near-linear deformations at smaller scales. Therefore, performing an additional more local linear registration would lead to a better fit. We subsequently crop square regions of interest (ROIs) with a size of 4,096 pixels ($\sim$5.3 mm) and without visible artifacts, covering distinct cellular architectures across the whole coronal plane. For all ROIs we perform additional affine registration and make sure that transformed landmarks have a maximum distance of 70 pixels (91 µm) from their matches. All ROIs are warped and resampled to 1.3 µm using linear interpolation to match the coordinate space of 3D-PLI. While this results in a loss of precision relative to the original resolution of Cresyl violet scans of 1 µm, matching the resolutions of both modalities facilitates subsequent processing and analysis steps.
@@ 2.3 Fourier-based online registration of image patches
To correct the remaining misalignment of 3D-PLI and Cresyl violet after affine registration of ROIs at a finer local scale, we introduce an online registration head that performs cross-modality alignment during training based on model predictions of small image patches. We assume that once the style transfer model has learned to reconstruct microscopic landmarks (e.g., individual cells or small blood vessels), such online registration will promote the learning of additional landmarks until the training can use pixel-aligned training examples.
The registration method performed during training needs to be computationally efficient, since training will require numerous registration iterations. Conventional feature-based image registration methods are accurate and can model nonlinear deformations but are computationally expensive and sensitive to image degradation. As we assume deformations to be approximately linear at a local scale, we take advantage of Fourier-based image correlations Tong et al., (2019), which can efficiently recover a translation between images in the frequency domain.
@@ 2.3.1 Translational shift
A common approach to retrieve the translational shift between $f$ and $g$ is to use circular cross-correlation Tong et al., (2019), which can be efficiently computed in the frequency domain as
@@
Eq. 2 can be extended to the mean over squared distances between pixel values as
@@
The translational shift $(\Delta u,\Delta v)$ can be recovered analog to Eq. 3 by computing the $\operatorname*{arg\,min}$. For periodic image functions $f$ and $g$, solutions of CC° and MSE° are identical as the sum over squared functions $f^{2}$ and $g^{2}$ is constant Fienup, (1997).
Since histological images are not periodic, zero padding is applied to fill both images up to a shape of $(H_{f}+H_{g}-1,W_{f}+W_{g}-1)$, in order to break periodicity and allow processing of images with different heights $H_{f}$, $H_{g}$ and widths $W_{f}$, $W_{g}$. The zero-padded images are denoted by new image functions $f_{0}$ and $g_{0}$. Furthermore, additional masks $M_{f}$ and $M_{g}$ are introduced, which have a value of one at all pixel coordinates within original height and width and zero elsewhere. We reformulate Eq. 4 to a non-circular form as
@@
which can be efficiently computed by multiple applications of Eq. 2. Here, the cross-correlations of $f_{0}^{2}$ and $g_{0}^{2}$ with masks $M_{g}$ and $M_{f}$, respectively, ensure that pixel values of the original unpadded images are not compared with zero padding values. Furthermore, the score is divided by the correlation between $M_{f}$ and $M_{g}$ to account for the number of overlapping pixels between original unpadded images. We apply the same division by $M_{f}\star M_{g}$ also for CC° to retrieve a non-circular variant called CC. While the solutions for circular CC° and MSE° are identical, the solutions for non-circular CC and MSE differ, resulting in distinct registration metrics.
@@ 2.3.2 Rotation and scale
While relative scale and rotation between images can also be retrieved in the frequency domain Sheng, (1989); Reddy and Chatterji, (1996), we expect only small relative rotation angles and minor scale variations due to the initial affine registration by matching blood vessels. Therefore, we leverage the parallel processing capability of GPUs to perform an exhaustive search over a fixed set of rotation angles without scaling adjustments. We calculate registration metrics for every translational shift and rotation and select the rotation angle and shift combination that yields a global optimum.
@@ 2.4 Conditional generation of Cresyl violet staining
Image-to-image translation refers to the process of generating images using a generator model $G$ conditioned on an input image $x$. Predictions $G(x)$ of the model are compared with actual target images $y$.
Our style transfer algorithm consists of three main elements:
  1. 1.
@@
    Reconstruction Loss $\mathcal{L}_{R}$ : Similar to Isola et al., (2017), we use an $L_{1}$ loss between target $y$ and prediction $G(x)$ to encourage pixel correspondence as
    $\mathcal{L}_{R}=\mathbb{E}_{x,y}\left\lVert R(y)-G(x)\right\rVert_{1},$ (6)
@@
    where $R$ performs the proposed online registration to spatially align $y$ and $G(x)$ before loss calculation. This enables the utilization of imperfectly aligned training data to penalize any discrepancies in corresponding pixel values between $G(x)$ and $y$.
  2. 2.
@@
    Style Loss $\mathcal{L}_{S}$: We apply a texture-sensitive style loss proposed by Gatys et al., (2015) based on squared distances between Gram matrix representations of neural network features. Given a pre-trained VGG encoder Simonyan and Zisserman, (2015), Gram matrix representations are computed from feature activations of its layers to characterize the texture of images at different complexities. For each layer $l$, the encoder produces a different number of $N_{l}$ feature maps, each storing $K_{l}$ spatial entries (i.e., height $\times$ width). Elements of the Gram matrix $G^{l}_{ij}$ at layer $l$ are computed as the inner product between the $i$-th and $j$-th feature map $F_{i}^{l}$ and $F_{j}^{l}$, where each map is flattened to a $K_{l}$-dimensional vector:
    $G^{l}_{ij}=\sum^{K_{l}}_{k=1}F^{l}_{ik}F^{l}_{jk}.$ (7)
@@
    Since the Gram matrix computation captures global feature correlations rather than spatial locations, this allows a translation-invariant comparison of image statistics. The style loss is computed over all $L$ layers of the VGG encoder, using Gram matrix representations $G^{l}$ for the online registered target and $\hat{G}^{l}$ for the prediction:
    $\mathcal{L}_{S}=\mathbb{E}_{x,y}\left[\sum^{L}_{l=1}\frac{1}{K_{l}^{2}N_{l}^{2}}\sum^{N_{l}}_{i=1}\sum^{N_{l}}_{j=1}(G^{l}_{ij}-\hat{G}^{l}_{ij})^{2}\right].$ (8)
    We refer to this style loss as Gram loss.
  3. 3.
@@
    Equivariance Loss $\mathcal{L}_{E}$: By registering target $y$ to generator prediction $G(x)$ before loss calculation, displacing objects in $G(x)$ relative to its input $x$ is not penalized by reconstruction loss $\mathcal{L}_{R}$. To prevent pixel shifts in $G(x)$, we introduce a third loss component enforcing equivariance with respect to rotations as
    $\mathcal{L}_{E}=\mathbb{E}_{x}\left\lVert\Omega(G(x))-G(\Omega(x))\right\rVert_{2},$ (9)
@@
    where operator $\Omega$ represents an image rotation of 180°180°180  °. Minimizing Eq. 9 ensures that pixels in $G(x)$ correspond to pixels at the same pixel coordinates in $x$ as any discrepancy would cause a mismatch of $\Omega(G(x))$ and $G(\Omega(x))$.
All components are aggregated into total loss
@@
with relative weightings $\lambda\in[0,1]$ and $\eta\geq 0$ as hyperparameters. An overview of the framework is illustrated in Fig. 3.
As an alternative for style loss $\mathcal{L}_{S}$, we consider adversarial training Goodfellow et al., (2014) using a GAN loss to compare with previous work in virtual staining Rivenson et al., (2019); Cross-Zamirski et al., (2022). We implement the GAN loss in the form of a Wasserstein GAN Arjovsky et al., (2017). In contrast to conditional GAN training Mirza and Osindero, (2014); Isola et al., (2017), we do not condition the discriminator on input images $x$, as this would cause the model to reproduce any misalignment in the training data. We refer to models trained with Gram loss for $\mathcal{L}_{S}$ as Gram and models trained with GAN loss as GAN. In combination with the online registration head, we refer to the models as Gram+Reg and GAN+Reg.
@@ 2.5 Model training
For the generator $G$, we use the same 5-layer U-Net Ronneberger et al., (2015) with numbers of features [32, 64, 128, 256, 512] in all experiments, and adjust the input and output channels to 3 according to our setup. To train $G$, we use square 3D-PLI patches of 444 pixels size, represented by parameter maps transmittance ($I_{T}$), retardation ($\sin\delta$) and direction ($\varphi$). We reformulate the 3D-PLI parameters as triplets $(I_{T},\sin\delta\cos(2\varphi),\sin\delta\sin(2\varphi))$ to resolve the circular behavior of direction $\varphi$, standardize the channels and stack them to the input of generator $G$. Due to the fully convolutional approach of the U-Net model without padding, the generated output predictions have a reduced size of 260 pixels. Unless specified otherwise, we use a patch size of 360 pixels for the target Cresyl violet images, centered at the input patch position. They are chosen to be larger than the model predictions to allow the online registration to correct translational shifts of up to 50 pixels in any direction, while keeping the predictions fully contained within the target images. We normalize image pixel values of the Cresyl violet staining to the range of [0, 1].
For computing style loss $\mathcal{L}_{S}$, we extract features from a VGG19 model Simonyan and Zisserman, (2015) to compute the Gram loss. The VGG feature encoder network has a depth of four layers and three input channels with pre-trained weights on ImageNet Deng et al., (2009). We multiply the style loss $\mathcal{L}_{S}$ by a constant factor of 104 to bring it to the same order of magnitude as reconstruction loss $\mathcal{L}_{R}$. For the training, we use Adam optimizer Kingma and Ba, (2017) with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999 and a learning rate of 10-3. If not stated otherwise, we use $\eta$ = 0.1 and $\lambda$ = 0.5 as default in the training objective in Eq. 10.
In the case of GAN style loss $\mathcal{L}_{S}$, we use a 4-layer convolutional network as discriminator with kernel size 4, stride 2, padding 1, and feature size of [32, 64, 128, 256], followed by batch normalization Ioffe and Szegedy, (2015) and Leaky ReLU after each convolution. For training, we use the Wasserstein GAN Arjovsky et al., (2017) objective and a separate Adam optimizer for the discriminator and the generator, using $\beta_{1}$ = 0.5, $\beta_{2}$ = 0.999 and a learning rate of 10-4. We perform five updates for the discriminator for one update of the generator and clamp discriminator weights at 0.03 after each step.
@@
For model training and evaluation, we use eight coronal sections at the level of the central sulcus. Seven sections are used for training and one section is kept for testing with a gap of 0.6 mm between train and test sections (Fig. 4 A). From the training sections, 27 affine-aligned ROIs are extracted (Fig. 4 B), where one ROI is held out for validation to identify possible overfitting. For each ROI, we extract joint target Cresyl violet images and 3D-PLI modalities at the same center location. To maximize the diversity of the training examples, we do not pre-compute training patches but sample them randomly during the training process. We use a batch size of 128 and draw 32,768 paired random patches per epoch evenly distributed across the training ROIs, resulting in approximately 1,260 random samples per ROI per epoch. All training is performed for 150 epochs or until model convergence if validation loss did not decrease for at least 50 epochs.
@@
To enhance the robustness of our trained models we employ 3D-PLI-specific data augmentations, which were carefully modified to keep physically plausible signal parameters Oberstrass et al., (2024). Specifically, we perform random rotation by angles between -180° and 180° with mirror padding and horizontal and vertical random flipping. In both cases, direction parameter maps $\varphi$ are corrected accordingly. Additionally, we perform Gaussian blurring of 3D-PLI parameter maps for random standard deviations up to $\sigma$ =1.5 and kernel sizes of 3 or 5. We scale thickness and attenuation coefficients for 3D-PLI parameter maps by random values between 0.5 and 2.
@@
All models were trained on the supercomputer JURECA-DC at the Jülich Supercomputing Centre (JSC, Thörnig, 2021) on a single node by splitting each batch equally onto 4 NVIDIA A100 GPUs using distributed data-parallel strategy. For data pre-processing 128 worker processes were spawned on 128 CPU cores. For reference, training for 100 epochs took 8 hours with online registration and 4 hours without on this hardware. The implementations are based on the Quicksetup-ai template by the HelmholtzAI Consultants Munich [^1] https://github.com/HelmholtzAI-Consultants-Munich/Quicksetup-ai, using the frameworks: PyTorch Paszke et al., (2019), PyTorch Lightning Borovec et al., (2022) and Hydra Yadan, (2019).
@@
For the online registration head, we restrict accepted solutions of Eq. 5 to translation + rotation pairs that cause the registered target to have full overlap with the prediction, avoiding loss calculation over zero-padded values. We check the translation correction for 31 rotation angles from -7.5° to 7.5° with steps of 0.5° and take the translation + rotation pair with the best registration score.

### v2

@@ 2.1.3 Tissue shrinkage estimation
We first estimate shrinkage in the 3D-PLI acquisition by affine registration of the 3D reconstructed blockface volume to the MRI. The linear part of the affine transformation in physical space has eigenvalues [0.984, 1.005, 1.015], indicating a global volume change of less than 1% and no axis-specific systematic deviation. In a second step, we calculate 2D shrinkage factors for each brain section as the quotient between the area occupied by tissue in the 3D-PLI measurement and its corresponding blockface image. For test section #559, we estimate a global 2D shrinkage factor of 0.97, indicating a slight swelling of the 3D-PLI measurement relative to its original area within the MRI. This observation is consistent across all sections, with 2D shrinkage factors between 0.95 and 0.99. We correct the in-plane sizes of segmented cell body areas in 3D-PLI by applying the individual 2D shrinkage factor of each section, assuming an approximately uniform area change of cells and surrounding tissue.
@@ 2.2 Initial cross-modality alignment
Performing the initial affine registration reveals remaining nonlinear deformations as shown in Fig.˜2D. Since nonlinear deformations typically have low spatial frequencies, causing smooth, large-scale distortions, we expected near-linear deformations at smaller scales. Therefore, performing an additional more local linear registration would lead to a better fit. We subsequently crop square regions of interest (ROIs) with a size of 4,096 pixels ($\sim$5.3 mm) and without visible artifacts, covering distinct cellular architectures across the whole coronal plane. For all ROIs we perform additional affine registration and make sure that transformed landmarks have a maximum distance of 70 pixels (91 µm) from their matches (Fig.˜1A). All ROIs are warped and resampled to 1.3 µm using linear interpolation to match the coordinate space of 3D-PLI. While this results in a loss of precision relative to the original resolution of Cresyl violet scans of 1 µm, matching the resolutions of both modalities facilitates subsequent processing and analysis steps.
@@ 2.3 Fourier-based online registration of image patches
To correct the remaining misalignment of 3D-PLI and Cresyl violet after affine registration of ROIs at a finer local scale, we introduce an online registration head that performs cross-modality alignment during training based on model predictions of small image patches (Fig.˜1B). We assume that once the style transfer model has learned to reconstruct microscopic landmarks (e.g., individual cells or small blood vessels), such online registration will promote the learning of additional landmarks until the training can use pixel-aligned training examples.
The registration method performed during training needs to be computationally efficient, since training will require numerous registration iterations. Conventional feature-based image registration methods are accurate and can model nonlinear deformations but are computationally expensive and sensitive to image degradation. As we assume deformations to be approximately linear at a local scale, we take advantage of Fourier-based image correlations tong2019, which can efficiently recover a translation between images in the frequency domain.
@@ 2.3.1 Translational shift
A common approach to retrieve the translational shift between $f$ and $g$ is to use circular cross-correlation tong2019, which can be efficiently computed in the frequency domain as
@@
Eq.˜2 can be extended to the mean over squared distances between pixel values as
@@
The translational shift $(\Delta u,\Delta v)$ can be recovered analog to Eq.˜3 by computing the $\operatorname*{arg\,min}$. For periodic image functions $f$ and $g$, solutions of CC° and MSE° are identical as the sum over squared functions $f^{2}$ and $g^{2}$ is constant fienup1997.
Since histological images are not periodic, zero padding is applied to fill both images up to a shape of $(H_{f}+H_{g}-1,W_{f}+W_{g}-1)$, in order to break periodicity and allow processing of images with different heights $H_{f}$, $H_{g}$ and widths $W_{f}$, $W_{g}$. The zero-padded images are denoted by new image functions $f_{0}$ and $g_{0}$. Furthermore, additional masks $M_{f}$ and $M_{g}$ are introduced, which have a value of one at all pixel coordinates within original height and width and zero elsewhere. We reformulate Eq.˜4 to a non-circular form as
@@
which can be efficiently computed by multiple applications of Eq.˜2. Here, the cross-correlations of $f_{0}^{2}$ and $g_{0}^{2}$ with masks $M_{g}$ and $M_{f}$, respectively, ensure that pixel values of the original unpadded images are not compared with zero padding values. Furthermore, the score is divided by the correlation between $M_{f}$ and $M_{g}$ to account for the number of overlapping pixels between original unpadded images. We apply the same division by $M_{f}\star M_{g}$ also for CC° to retrieve a non-circular variant called CC. While the solutions for circular CC° and MSE° are identical, the solutions for non-circular CC and MSE differ, resulting in distinct registration metrics.
@@ 2.3.2 Rotation and scale
While relative scale and rotation between images can also be retrieved in the frequency domain sheng1989; reddy1996, we expect only small relative rotation angles and minor scale variations due to the initial affine registration by matching blood vessels. Therefore, we leverage the parallel processing capability of GPUs to perform an exhaustive search over a fixed set of rotation angles without scaling adjustments. We calculate registration metrics for every translational shift and rotation and select the rotation angle and shift combination that yields a global optimum.
@@ 2.4 Conditional generation of Cresyl violet staining
Image-to-image translation refers to the process of generating images using a generator model $G$ conditioned on an input image $x$. Predictions $G(x)$ of the model are compared with actual target images $y$. In our case, we translate from the domain of 3D-PLI images to Cresyl violet. The translation is performed by a U-Net ronneberger2015 serving as the generator model, which forms the core of our image-to-image translation framework, as illustrated in Fig.˜3. The three loss components used to train the U-Net are described below.
2.4.1 Reconstruction Loss
@@
Similar to isola2017, we use an $L_{1}$ loss between target $y$ and prediction $G(x)$ to encourage pixel correspondence in the reconstruction loss
$\mathcal{L}_{R}=\mathbb{E}_{x,y}\left[\left\lVert R_{y,G(x)}(y)-G(x)\right\rVert_{1}\right],$ (6)
@@
where $R_{y,G(x)}$ performs the proposed online registration to spatially align $y$ and $G(x)$ before loss calculation. This enables utilization of imperfectly aligned training data to penalize any discrepancies in corresponding pixel values between $G(x)$ and $y$. To disable online registration, $R_{y,G(x)}$ can be replaced by the identity function.
2.4.2 Style Loss
We consider two alternative implementations of a style loss $\mathcal{L}_{S}$, both focussing on style preservation. We refer to the first one as Gram loss, and the second as GAN loss.
@@ 2.4.2 Style Loss
For the Gram loss, we apply a texture-sensitive style loss proposed by gatys2015a based on squared distances between Gram matrix representations of neural network features. Given a pre-trained VGG encoder simonyan2015, Gram matrix representations are computed from feature activations of its layers to characterize the texture of images at different complexities. For each layer $l$, the encoder produces a different number of $N_{l}$ feature maps, each storing $K_{l}$ spatial entries (i.e., height $\times$ width). Elements of the Gram matrix $\Gamma^{l}_{ij}$ at layer $l$ are computed as the inner product between the $i$-th and $j$-th feature map $F_{i}^{l}$ and $F_{j}^{l}$, where each map is flattened to a $K_{l}$-dimensional vector:
$\Gamma^{l}_{ij}=\sum^{K_{l}}_{k=1}F^{l}_{ik}F^{l}_{jk}.$ (7)
@@
Since the Gram matrix computation captures global feature correlations rather than spatial locations, this allows a translation-invariant comparison of image statistics. The style loss is computed over all $L$ layers of the VGG encoder, using Gram matrix representations $\Gamma^{l}$ for the online registered target and $\hat{\Gamma}^{l}$ for the prediction:
$\mathcal{L}_{S}=\mathbb{E}_{x,y}\left[\sum^{L}_{l=1}\frac{1}{K_{l}^{2}N_{l}^{2}}\sum^{N_{l}}_{i=1}\sum^{N_{l}}_{j=1}(\Gamma^{l}_{ij}-\hat{\Gamma}^{l}_{ij})^{2}\right].$ (8)
For the GAN loss, we consider adversarial training goodfellow2014 to compare with previous work in virtual staining rivenson2019; cross-zamirski2022. We implement the GAN loss in the form of a Wasserstein GAN arjovsky2017a. In contrast to conditional GAN training mirza2014; isola2017, we do not condition the discriminator on input images $x$, as this would cause the model to reproduce any misalignment in the training data.
2.4.3 Equivariance Loss
@@
By registering target $y$ to generator prediction $G(x)$ before loss calculation, displacements of objects in $G(x)$ relative to $x$ are not captured by $\mathcal{L}_{R}$. To prevent pixel shifts in $G(x)$, we enforce equivariance with respect to rotations through the equivariance loss
$\mathcal{L}_{E}=\mathbb{E}_{x}\left[\left\lVert\Omega(G(x))-G(\Omega(x))\right\rVert_{2}\right],$ (9)
@@
where operator $\Omega$ represents an image rotation of 180°. Minimizing Eq.˜9 ensures that pixels in $G(x)$ correspond to pixels at the same pixel coordinates in $x$ as any discrepancy would cause a mismatch of $\Omega(G(x))$ and $G(\Omega(x))$.
2.4.4 Total loss formulation
All components are aggregated into the total loss
@@ 2.4.4 Total loss formulation
with relative weightings $\lambda\in[0,1]$ and $\eta\geq 0$ as hyperparameters. We denote models that use Gram loss as the style loss $\mathcal{L}_{S}$ as Gram, and models that use GAN loss as the style loss $\mathcal{L}_{S}$ as GAN. When online registration is enabled during computation of reconstruction loss $\mathcal{L}_{R}$, the corresponding models are denoted as Gram+Reg and GAN+Reg. Base models Gram and GAN compute the reconstruction loss with online registration disabled.
@@ 2.5 Model training
For the generator $G$, we use the same 5-layer U-Net ronneberger2015 with numbers of features [32, 64, 128, 256, 512] in all experiments, and adjust the input and output channels to 3 according to our setup. To train $G$, we use square 3D-PLI patches of 444 pixels size, represented by parameter maps transmittance ($I_{T}$), retardation ($\sin\delta$) and direction ($\varphi$). We reformulate the 3D-PLI parameters as triplets $(I_{T},\sin\delta\cos(2\varphi),\sin\delta\sin(2\varphi))$ to resolve the circular behavior of direction $\varphi$, standardize the channels and stack them to the input of generator $G$. Due to the fully convolutional approach of the U-Net model without padding, the generated output predictions have a reduced size of 260 pixels. Unless specified otherwise, we use a patch size of 360 pixels for the target Cresyl violet images, centered at the input patch position. They are chosen to be larger than the model predictions to allow the online registration to correct translational shifts of up to 50 pixels in any direction, while keeping the predictions fully contained within the target images. We normalize image pixel values of the Cresyl violet staining to the range of [0, 1].
For computing style loss $\mathcal{L}_{S}$, we extract features from a VGG19 model simonyan2015 to compute the Gram loss. The VGG feature encoder network has a depth of four layers and three input channels with pre-trained weights on ImageNet deng2009. We multiply the style loss $\mathcal{L}_{S}$ by a constant factor of 104 to bring it to the same order of magnitude as reconstruction loss $\mathcal{L}_{R}$. For the training, we use Adam optimizer kingma2017 with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999 and a learning rate of 10-3. If not stated otherwise, we use $\eta$ = 0.1 and $\lambda$ = 0.5 as default in the training objective in Eq.˜10.
In the case of GAN style loss $\mathcal{L}_{S}$, we use a 4-layer convolutional network as discriminator with kernel size 4, stride 2, padding 1, and feature size of [32, 64, 128, 256], followed by batch normalization ioffe2015 and Leaky ReLU after each convolution. For training, we use the Wasserstein GAN arjovsky2017a objective and a separate Adam optimizer for the discriminator and the generator, using $\beta_{1}$ = 0.5, $\beta_{2}$ = 0.999 and a learning rate of 10-4. We perform five updates for the discriminator for one update of the generator and clamp discriminator weights at 0.03 after each step.
@@
For model training and evaluation, we use eight coronal sections at the level of the central sulcus. Seven sections are used for training and one section is kept for testing with a gap of 0.6 mm between train and test sections (Fig.˜4A). From the training sections, 27 affine-aligned ROIs are extracted (Fig.˜4B), where one ROI is held out for validation to identify possible overfitting. For each ROI, we extract joint target Cresyl violet images and 3D-PLI modalities at the same center location. To maximize the diversity of the training examples, we do not pre-compute training patches but sample them randomly during the training process. We use a batch size of 128 and draw 32,768 paired random patches per epoch evenly distributed across the training ROIs, resulting in approximately 1,260 random samples per ROI per epoch. All training is performed for 150 epochs or until model convergence if validation loss did not decrease for at least 50 epochs.
@@
To enhance the robustness of our trained models we employ 3D-PLI-specific data augmentations, which were carefully modified to keep physically plausible signal parameters oberstrass2024a. Specifically, we perform random rotation by angles between -180° and 180° with mirror padding and horizontal and vertical random flipping. In both cases, direction parameter maps $\varphi$ are corrected accordingly. Additionally, we perform Gaussian blurring of 3D-PLI parameter maps for random standard deviations up to $\sigma$ =1.5 and kernel sizes of 3 or 5. We scale thickness and attenuation coefficients for 3D-PLI parameter maps by random values between 0.5 and 2.
@@
All models were trained on the supercomputer JURECA-DC at the Jülich Supercomputing Centre (JSC, thornig2021) on a single node by splitting each batch equally onto 4 NVIDIA A100 GPUs using distributed data-parallel strategy. For data pre-processing 128 worker processes were spawned on 128 CPU cores. For reference, training for 100 epochs took 8 hours with online registration and 4 hours without on this hardware. The implementations are based on the Quicksetup-ai template by the HelmholtzAI Consultants Munich [^1] https://github.com/HelmholtzAI-Consultants-Munich/Quicksetup-ai, using the frameworks: PyTorch NEURIPS2019_9015, PyTorch Lightning jirka_borovec_2022_7447212 and Hydra Yadan2019Hydra.
@@
For the online registration head, we restrict accepted solutions of Eq.˜5 to translation + rotation pairs that cause the registered target to have full overlap with the prediction, avoiding loss calculation over zero-padded values. We check the translation correction for 31 rotation angles from -7.5° to 7.5° with steps of 0.5° and take the translation + rotation pair with the best registration score.